OpenStudy (anonymous):

Evaluate the series 4+2+1+1/2+1/4 to S10

OpenStudy (raden):

use the formula bellow : Sn = a1(1 - r^n)/(1 - r) with a1 = the first term r = ratio n = number terms (known n = 10)

OpenStudy (raden):

to get ratio : r = a2/a1 = 2/4 = 1/2

OpenStudy (anonymous):

how do you find the ratio

OpenStudy (raden):

look at my the 2nd coment ^

OpenStudy (anonymous):

yeah did not see that at first

OpenStudy (raden):

now, distribute all values above into that formula to get your answer

OpenStudy (anonymous):



OpenStudy (raden):

yeah, that can be corrected :) i got 7.9921875

OpenStudy (anonymous):

thanks

OpenStudy (raden):

you're welcome